### v1

@@ 6.3. The general case
If $Z\subset X$ and $Z^{\prime}\subset X^{\prime}$ are closed subvarieties, and if $f:X\to X^{\prime}$ is a morphism such that $Z=f^{-1}Z^{\prime}$, then there is a ${\mathbb{G}}_{m}$-equivariant morphism $F:V(Z/X)\to V(Z^{\prime}/X^{\prime})$ which induces a map of the corresponding diagrams above. If $f$ is proper, then $F$ is proper.
@@
Suppose $Z=f^{-1}Z^{\prime}$, then $f$ induces a map $F_{1}:B(Z/X)\to B(Z^{\prime}/X^{\prime})$ between the blow ups such that $F_{1}^{-1}(Bl_{Z^{\prime}}X^{\prime}\times\{0\})=Bl_{Z}X\times\{0\}$. This restricts to give $F:V(Z/X)\to V(Z^{\prime}/X^{\prime})$. If $f$ is proper, then $F_{1}$ is proper, so the same is true for $F$. ∎
@@
Suppose we are given an admissible datum $(X,Y,f,U,Z)$ such that $Y$ and $Z$ have the $PEC$ property. Since $Y$ is quasiprojective, we can choose a Cartier divisor $D\supseteq f(Z)$. Let us write $f^{-1}D=Z\cup Z^{\prime}$, where $Z^{\prime}$ is the union of components $Z_{i}^{\prime}$ not contained in $Z$. Then $(Z_{i}^{\prime},f(Z_{i}^{\prime}),f|_{Z_{i}^{\prime}},U\cap Z_{i}^{\prime},Z_{i}^{\prime}\cap Z)$ are admissible data satisfying the assumptions of the theorem. By induction and Lemma 6.5, we find that $Z^{\prime}$ has the $PEC$ property. Since $Z$ also has the $PEC$ property, we can conclude that so does $Z\cup Z^{\prime}$ by Lemma 6.5. Therefore $C(Z\cup Z^{\prime}/X)$ has the $PEC_{m}[-1]$ property by Lemma 6.8. This Lemma also shows that $V_{Y}:=V(D/Y)$ has the $PEC_{m}[-1]$-property. Let $V_{X}=V(Z\cup Z^{\prime}/X)$. The map $f$ induces a proper map $F:V_{X}\to V_{Y}$ by Lemma 6.7. Therefore we have admissible datum $(V_{X},V_{Y},F,U\times{\mathbb{G}}_{m},Z^{\prime\prime})$, where $Z^{\prime\prime}=C(Z\cup Z^{\prime}/X)\cup(Z\cup Z^{\prime})\times{\mathbb{A}}^{1}$. Lemma 6.5 shows that $Z^{\prime\prime}$ has the $PEC_{m}[-1]$ property. Moreover, we have $h^{-1}(0)\supseteq F(Z^{\prime\prime})$, where $h:V_{Y}\rightarrow{\mathbb{A}}^{1}$ is the projection. In particular, we are in the setting of Lemma 6.6. It follows that $V_{X}$ has the $PEC_{m}[-1]$-property, and therefore $X$ has the $PEC$-property by Lemma 6.8.
@@ 7. Extended Torelli map
Let $\overline{{\mathscr{M}}_{g}}$ (respectively ${\mathscr{M}}_{g}$) denote the stack parameterizing stable (respectively smooth) curves of genus $g$ [9]. Recall that a stable curve $C$ is of compact type if its dual graph is a tree, or equivalently if the identity component of its Picard scheme $\operatorname{Pic}^{0}(C)$ is an abelian variety. Let
@@
By a theorem of Mumford and Namikawa [26, §9D] the Torelli map $\tau:{\mathscr{M}}_{g}\to{\mathscr{A}}_{g}$ extends to a map $\tau:\overline{{\mathscr{M}}_{g}}\to\overline{{\mathscr{A}}_{g}}$ to the toroidal compactification associated to the second Voronoi decomposition. Alexeev [2, 3] gave a modular interpretation of both the target and the map. On ${\mathscr{M}}_{g}^{c}$, $\tau$ simplifies to
[…]
@@
Let $\overline{{\mathscr{M}}_{g,n}}$ be the Deligne-Mumford stack of genus $g$ stable curves with $n$ base points [19]. A geometric point of this space stack corresponds a semistable curve $(C,p_{1},\ldots,p_{n})$ of arithmetic genus $g$ with $n$ smooth base points such that the total number of base and double points on any smooth rational component is at least three. We refer to this is as an $n$-pointed stable curve. When $n>0$, there exists a morphism $\pi_{n}:\overline{{\mathscr{M}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n-1}}$ called a contraction. For a curve without rational components, $\pi_{n}$ simply forgets $p_{n}$. Given integers $g=g_{1}+g_{2}$, and a partition of $\{1,\ldots,n\}$ into two subsets of cardinality $n_{1}$ and $n_{2}$, Knudsen also constructs a gluing map
@@
We recall two basic facts regarding these constructions (see [19]).
@@
    First, the contraction map is equivalent to the universal curve over ${\mathscr{M}}_{g,n-1}$ (see Lemma 7.2).
@@
The forgetful map $\overline{{\mathscr{M}}_{g,n}}\to\overline{{\mathscr{M}}_{g}}$ is the composition of contractions. Let ${{\mathscr{M}}_{g,n}}$ and ${{\mathscr{M}}^{c}_{g,n}}$ denote the preimages of ${\mathscr{M}}_{g}$ and ${\mathscr{M}}_{g}^{c}$ respectively. We note that ${\mathscr{M}}^{c}_{1,n}={\mathscr{M}}_{1,n}$. When $g>1$, the complement ${\mathscr{M}}^{c}_{g,n}-{\mathscr{M}}_{g,n}$ is a union of divisors, each of which can be described as the image of a gluing map with respect to a partition $g=g_{1}+g_{2}$ with $g_{i}>0$.
@@
Given a scheme $S$, an $S$-valued point of $\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{M}}^{c}_{g,n+1}$) is a stable curve ${\mathcal{C}}\to S$ with $n$ sections $\sigma_{1},\ldots,\sigma_{n}$, such that the geometric fibres are $n$-pointed genus $g$ stable curves (of compact type). The universal curve $\overline{{\mathscr{Z}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{Z}}^{c}_{g,n}\to{\mathscr{M}}^{c}_{g,n}$) is the stack whose $S$-points are stable curves ${\mathcal{C}}\to S$ with $n+1$ sections $\sigma_{1},\ldots,\sigma_{n+1}$, such that $({\mathcal{C}}\to S,\sigma_{1},\ldots,\sigma_{n})$ defines a point of $\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{M}}^{c}_{g,n}$). There is an obvious functor $\overline{{\mathscr{M}}_{g,n+1}}\to\overline{{\mathscr{Z}}_{g,n}}$, which sends to ${\mathscr{M}}^{c}_{g,n+1}$ to ${\mathscr{Z}}^{c}_{g,n}$. Knudsen [19, §2] defines an inverse $s:\overline{{\mathscr{Z}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n+1}}$ called stabilization. When $S=\operatorname{Spec}k$, then roughly
$s({\mathscr{C}},\sigma_{1},\ldots,\sigma_{n+1})=\begin{cases}({\mathcal{C}},\sigma_{1},\ldots,\sigma_{n+1})&\text{if this is an $(n+1)$-pointed stable curve}\\ ({\mathcal{C}}^{\prime},\sigma_{1}^{\prime},\ldots,\sigma_{n+1}^{\prime})&\text{if $\sigma_{n+1}$ is a singular point or another section}\\ \end{cases}$
@@
For existence, it suffices to treat the case where $N\geq 3$, since the other cases can be handled by taking quotients. Then $M={}_{N}{\mathscr{M}}^{c}_{g,n}$ is fine with a universal family ${\mathcal{C}}={}_{N}{\mathscr{Z}}_{g,n}$ over it. The curve ${\mathcal{C}}$ carries universal sections $\sigma_{1},\ldots,\sigma_{n}$. The relative Picard algebraic space $\operatorname{Pic}({\mathcal{C}}/M)$ exists by [8, chap 9.4, thm 1]. Furthermore, this is a disjoint union of subschemes $P^{d}=\operatorname{Pic}^{d}({\mathcal{C}}/M)$ parameterizing line bundles with degree (defined as $\chi(L)-g+1$) $d$ on the fibres. Note that $P^{0}\to M$ is a $g$-dimensional abelian scheme with level $N$-structure, and $P^{d}$ is a torsor over $P^{0}$. When $d=2g-2$, there is a canonical section given by the class of $\omega_{{\mathcal{C}}/M}$, which allows us to identify $P^{2g-2}$ with $P^{0}$. The class of $(2g-2)\sigma_{i}$ defines a section of $\operatorname{Pic}({\mathcal{C}}/M)$. In fact, by restricting to ${}_{N}{\mathscr{M}}_{g,n}$, we can see that $(2g-2)\sigma_{i}$ lies in $P^{2g-2}$. We can view these as sections of $P^{0}$ by the above identification. Therefore, the universal property of ${}_{N}{\mathscr{A}}_{g,n}$ gives map $\tau:M\to{}_{N}{{\mathscr{A}}_{g,n}}$. On points of $M$, it is clearly described (with some abuse of notation) as
@@
To check properness, we use the valuative criterion. Let $R$ be a discrete valuation ring with generic point $\eta$, and suppose that we are given a $2$-commutative diagram
@@
with solid arrows. Then after replacing $R$ by a finite extension, we will construct an arrow $h$ making the new diagram $2$-commute. Since $\overline{{\mathscr{M}}_{g,n}}$ is proper [19], we do get an arrow $f:\operatorname{Spec}R\to\overline{{\mathscr{M}}_{g,n}}$ (after a finite extension), making the resulting diagram commute. In geometric language, we are given a $g$-dimensional principally polarized abelian scheme ${\mathcal{A}}\to\operatorname{Spec}R$, with $n$ sections $\alpha_{i}$. The map $f$ gives a stable $n$-pointed genus $g$ curve $({\mathcal{C}}\to\operatorname{Spec}R,\sigma_{i})$ such that
@@
By the theory of Néron models [8, chap 9.5, thm 4] , we have ${\mathcal{A}}\cong\operatorname{Pic}^{0}({\mathcal{C}}/\operatorname{Spec}R)$. It follows that ${\mathcal{C}}$ is of compact type. Since ${\mathcal{A}}/\operatorname{Spec}R$ is separated, we have an isomorphism
@@
Therefore the image of $f$ lies in ${\mathscr{M}}^{c}_{g,n}$, and moreover $h=f$ makes the diagram commute. The extension is to the level $N$ case is straightforward and left to the reader. ∎
@@ 8. Euler characteristics of perverse sheaves on ${\mathscr{M}}_{g,n}$
If $P$ is a perverse sheaf on ${\mathscr{M}}_{g}$, then $\chi(j_{!*}P)\geq 0$.
@@
If $g=3$ then $\tau$ is semismall. Therefore $\chi({\mathscr{M}}_{3}^{c})\geq 0$
@@
In case (a), $C_{1}$ can be taken to be an elliptic curve with $p$ as its origin. Since $\tau([C])=[\operatorname{Pic}^{0}(C_{1})\times\operatorname{Pic}^{0}(C_{2})]$, the map forgets $p\in C_{2}$. Therefore the fibre $\tau^{-1}(\tau([C]))$ can be identified with $C_{2}$. In case (b), we take $C_{1}\cap C_{2}$ as the origins of $C_{1}$, $C_{2}$, and $C_{2}\cap C_{3}$ as the origin of $C_{3}$. The point $p\in C_{2}\cap C_{3}$ is distinct from 00 in this case; we treat (c) as the limiting case where $p\to 0$. Then $\tau^{-1}(\tau([C]))$ with $C_{2}$. The claim follows from this, and it implies that $\tau$ is also semismall, when $g=3$. The inequality follows from corollary 5.5. ∎
@@
We will prove the corresponding inequality for all $g$. However, the map ${\mathscr{M}}_{g}^{c}\to{\mathscr{A}}_{g}$ is not semismall when $g>3$, so it requires a different strategy.
@@
Let ${\mathscr{M}}={\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{c}$, where either $g_{i}>1$ or $g_{i}=1$ and $n_{i}>0$. If ${\mathcal{K}}$ is perverse sheaf on ${\mathscr{M}}$, then

### v2

@@ 5.5. The general case
If $Z\subset X$ and $Z^{\prime}\subset X^{\prime}$ are closed subvarieties, and if $f:X\to X^{\prime}$ is a morphism such that $Z=f^{-1}Z^{\prime}$, then there is a ${\mathbb{G}}_{m}$-equivariant morphism $F:V(Z/X)\to V(Z^{\prime}/X^{\prime})$ which induces a map of the corresponding diagrams above.
@@
Suppose $Z=f^{-1}Z^{\prime}$, then $f$ induces a map $F_{1}:B(Z/X)\to B(Z^{\prime}/X^{\prime})$ between the blow ups such that $F_{1}^{-1}(Bl_{Z^{\prime}}X^{\prime}\times\{0\})=Bl_{Z}X\times\{0\}$. This restricts to give $F:V(Z/X)\to V(Z^{\prime}/X^{\prime})$. ∎
@@
Suppose we are given an admissible datum $(X,Y,f,U,Z)$ such that $Y$ and $Z$ have the $PEC$ property. Since $Y$ is quasiprojective, we can choose a Cartier divisor $D\supseteq f(Z)$. Let us write $f^{-1}D=Z\cup Z^{\prime}$, where $Z^{\prime}$ is the union of components $Z_{i}^{\prime}$ not contained in $Z$. Then $(Z_{i}^{\prime},f(Z_{i}^{\prime}),f|_{Z_{i}^{\prime}},U\cap Z_{i}^{\prime},Z_{i}^{\prime}\cap Z)$ are admissible data satisfying the assumptions of the theorem. By induction and Lemma 5.13, we find that $Z^{\prime}$ has the $PEC$ property. Since $Z$ also has the $PEC$ property, we can conclude that so does $Z\cup Z^{\prime}$ by Lemma 5.13. Therefore $C(Z\cup Z^{\prime}/X)$ has the $PEC_{m}[-1]$ property by Lemma 5.16. This Lemma also shows that $V_{Y}:=V(D/Y)$ has the $PEC_{m}[-1]$-property. Let $V_{X}=V(Z\cup Z^{\prime}/X)$. The map $f$ induces a morphism $F:V_{X}\to V_{Y}$ by Lemma 5.15. Therefore we have admissible datum $(V_{X},V_{Y},F,U\times{\mathbb{G}}_{m},Z^{\prime\prime})$, where $Z^{\prime\prime}=C(Z\cup Z^{\prime}/X)\cup(Z\cup Z^{\prime})\times{\mathbb{A}}^{1}$. Lemma 5.13 shows that $Z^{\prime\prime}$ has the $PEC_{m}[-1]$ property. Moreover, we have $g^{-1}(0)\supseteq F(Z^{\prime\prime})$, where $g:V_{Y}\rightarrow{\mathbb{A}}^{1}$ is the projection. In particular, we are in the setting of Lemma 5.14. It follows that $V_{X}$ has the $PEC_{m}[-1]$-property, and therefore $X$ has the $PEC$-property by Lemma 5.16.
@@ 6. Extended Torelli map
Let $\overline{{\mathscr{M}}_{g}}$ (respectively ${\mathscr{M}}_{g}$) denote the stack parameterizing stable (respectively smooth) curves of genus $g$ [dm]. Recall that a stable curve $C$ is of compact type if its dual graph is a tree, or equivalently if the identity component of its Picard scheme $\operatorname{Pic}^{0}(C)$ is an abelian variety. Let
@@
By a theorem of Mumford and Namikawa [namikawa, §9D] the Torelli map $\tau:{\mathscr{M}}_{g}\to{\mathscr{A}}_{g}$ extends to a map $\tau:\overline{{\mathscr{M}}_{g}}\to\overline{{\mathscr{A}}_{g}}$ to the toroidal compactification associated to the second Voronoi decomposition. Alexeev [alexeev1, alexeev2] gave a modular interpretation of both the target and the map. On ${\mathscr{M}}_{g}^{c}$, $\tau$ simplifies to
$\tau([C])=[\operatorname{Pic}^{0}(C)]=[\prod_{i}(\operatorname{Pic}^{0}(C_{i})]\quad\text{(as principally polarized abelian varieties})$
@@
Let $\overline{{\mathscr{M}}_{g,n}}$ be the Deligne-Mumford stack of genus $g$ stable curves with $n$ base points [knudsen]. A geometric point of this stack corresponds a semistable curve $(C,p_{1},\ldots,p_{n})$ of arithmetic genus $g$ with $n$ smooth base points such that the total number of base and double points on any smooth rational component is at least three. We refer to this is as an $n$-pointed stable curve. When $n>0$, there exists a morphism $\pi_{n}:\overline{{\mathscr{M}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n-1}}$ called a contraction. For a curve without rational components, $\pi_{n}$ simply forgets $p_{n}$. Given integers $g=g_{1}+g_{2}$, and a partition of $\{1,\ldots,n\}$ into two subsets of cardinality $n_{1}$ and $n_{2}$, Knudsen also constructs a gluing map
@@
We recall two basic facts regarding these constructions (see [knudsen]).
@@
    First, the contraction map is equivalent to the universal curve over ${\mathscr{M}}_{g,n-1}$ (see Lemma 6.2). By equivalent we mean there is an isomorphism of stacks over ${\mathscr{M}}_{g,n-1}$.
@@
The forgetful map $\overline{{\mathscr{M}}_{g,n}}\to\overline{{\mathscr{M}}_{g}}$ is the composition of contractions. Let ${\mathscr{M}}_{g,n}$} and ${\mathscr{M}}^{c}_{g,n}$} denote the preimages of ${\mathscr{M}}_{g}$ and ${\mathscr{M}}_{g}^{c}$ respectively. The complement ${\mathscr{M}}^{c}_{g,n}-{\mathscr{M}}_{g,n}$ is a union of divisors, each of which can be described as the image of an aforementioned gluing map (intersected with ${\mathscr{M}}_{g,n}^{c}$).
@@
Given a scheme $S$, an $S$-valued point of $\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{M}}^{c}_{g,n+1}$) is a stable curve ${\mathcal{C}}\to S$ with $n$ sections $\sigma_{1},\ldots,\sigma_{n}$, such that the geometric fibres are $n$-pointed genus $g$ stable curves (of compact type). The universal curve $\overline{{\mathscr{Z}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{Z}}^{c}_{g,n}\to{\mathscr{M}}^{c}_{g,n}$) is the stack whose $S$-points are stable curves ${\mathcal{C}}\to S$ with $n+1$ sections $\sigma_{1},\ldots,\sigma_{n+1}$, such that $({\mathcal{C}}\to S,\sigma_{1},\ldots,\sigma_{n})$ defines a point of $\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{M}}^{c}_{g,n}$). There is an obvious functor $\overline{{\mathscr{M}}_{g,n+1}}\to\overline{{\mathscr{Z}}_{g,n}}$, which sends to ${\mathscr{M}}^{c}_{g,n+1}$ to ${\mathscr{Z}}^{c}_{g,n}$. Knudsen [knudsen, §2] defines an inverse $s:\overline{{\mathscr{Z}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n+1}}$ called stabilization. When $S=\operatorname{Spec}k$, then roughly
[…]
@@
For existence, it suffices to treat the case where $N\geq 3$, since the other cases can be handled by taking quotients. Then $M={}_{N}{\mathscr{M}}^{c}_{g,n}$ is fine with a universal family ${\mathcal{C}}={}_{N}{\mathscr{Z}}_{g,n}$ over it. The curve $\mathcal{C}$} carries universal sections $\sigma_{1},\ldots,\sigma_{n}$. The relative Picard algebraic space $\operatorname{Pic}({\mathcal{C}}/M)$ exists by [blr, chap 9.4, thm 1]. Furthermore, this is a disjoint union of subschemes $P^{d}=\operatorname{Pic}^{d}({\mathcal{C}}/M)$ parameterizing line bundles $\mathcal{L}$} of degree $d$ on the fibers, where we define the the degree to be $\chi(L)-g+1$. Note that $P^{0}\to M$ is a $g$-dimensional abelian scheme with level $N$-structure, and $P^{d}$ is a torsor over $P^{0}$. When d=2g2d=2g-2, there is a canonical section given by the class of $\omega_{{\mathcal{C}}/M}$, which allows us to identify $P^{2g-2}$ with $P^{0}$. The class of $(2g-2)\sigma_{i}$ defines a section of $\operatorname{Pic}({\mathcal{C}}/M)$. In fact, by restricting to $_{N}{\mathscr{M}}_{g,n}$, we can see that $(2g-2)\sigma_{i}$ lies in $P^{2g-2}$. We can view these as sections of $P^{0}$ by the above identification. Therefore, the universal property of $_{N}{\mathscr{A}}_{g,n}$ gives map $\tau:M\to{}_{N}{{\mathscr{A}}_{g,n}}$. On points of $M$, it is clearly described (with some abuse of notation) as
@@
To check properness, we use the valuative criterion for stacks ([dm, thm 4.19], or [stacks-project], Tag 0CLY, Tag 0CLK , Tag 0CLG). Let $R$ be a discrete valuation ring with generic point $\eta$, and suppose that we are given a $2$-commutative diagram
@@
with solid arrows. Then after replacing $R$ by a finite extension, we will construct an arrow $h$ making the new diagram $2$-commute. [^2] If $\eta=\operatorname{Spec}(K)$, then ‘after a finite extension of $R$’ means we pass to a finite extension $K^{\prime}$ of $K$ and a valuation ring $R^{\prime}\subset K^{\prime}$ dominating $R$. Since $\overline{{\mathscr{M}}_{g,n}}$ is proper [knudsen], we do get an arrow $f:\operatorname{Spec}R\to\overline{{\mathscr{M}}_{g,n}}$ (after a finite extension), making the resulting diagram commute. In geometric language, we are given a $g$-dimensional principally polarized abelian scheme ${\mathcal{A}}\to\operatorname{Spec}R$, with $n$ sections $\alpha_{i}$. The map $f$ gives a stable $n$-pointed genus $g$ curve $({\mathcal{C}}\to\operatorname{Spec}R,\sigma_{i})$ such that
@@
By the theory of Néron models [blr, chap 9.5, thm 4] , we have ${\mathcal{A}}\cong\operatorname{Pic}^{0}({\mathcal{C}}/\operatorname{Spec}R)$. It follows that $\mathcal{C}$} is of compact type. Since ${\mathcal{A}}/\operatorname{Spec}R$ is separated, we have an isomorphism
@@
Therefore the image of $f$ lies in ${\mathscr{M}}^{c}_{g,n}$, and moreover $h=f$ makes the diagram commute. The extension to the level $N$ case is straightforward and left to the reader. ∎
Let us say a pointed curve, corresponding to a point of ${\mathscr{M}}^{c}_{g,n}$, is of special compact type if the dual graph has the property that no two genus zero vertices are joined by an edge. In particular, the dual graph has no rational trees. We now let ${\mathscr{M}}_{g,n}^{sc}\subset{\mathscr{M}}_{g,n}^{c}$ denote the substack of curves of special compact type. This can be seen to be an open substack.
Note that ${\mathscr{M}}_{0,n}^{sc}={\mathscr{M}}_{0,n}$, and ${\mathscr{M}}_{1,1}^{sc}={\mathscr{M}}_{1,1}$. In general, the boundary divisors ${\mathscr{M}}_{g_{1},n_{1}+1}^{c}\times{\mathscr{M}}_{g_{2},n_{1}+1}^{c}$ of ${\mathscr{M}}^{c}_{g,n}$ restricted to ${\mathscr{M}}_{g,n}^{sc}$ give the boundary divisors ${\mathscr{M}}_{g_{1},n_{1}+1}^{sc}\times{\mathscr{M}}_{g_{2},n_{2}+1}^{sc}$ (where we have $g_{1}+g_{2}=g$ and $n_{1}+n_{2}=n$). For example, the boundary of ${\mathscr{M}}_{1,2}^{sc}$ is the divisor ${\mathscr{M}}_{1,1}\times{\mathscr{M}}_{0,3}$. We may consider analogs with the level $N$-structure ($g>0$) defined as before.
Lemma 6.4.
We claim that $\tau({\mathscr{M}}_{g,n}^{sc})=\tau({\mathscr{M}}_{g,n}^{c})$, and therefore this set is closed in ${\mathscr{A}}_{g,n}$ (since $\tau$ is proper on ${\mathscr{M}}_{g,n}^{c}$).
Proof.
We may check the claim on geometric points. If we have a curve with a rational chain in the dual graph, we can contract it to a single $\mathbb{P}^{1}$ attached to a higher genus curve or two different higher genus curves. This object is in ${\mathscr{M}}_{g,n}^{sc}$ and has the same image under $\tau$.
@@ 7. Euler characteristics of perverse sheaves on ${\mathscr{M}}_{g,n}$
If $P$ is a perverse sheaf on ${\mathscr{M}}_{g}$, then $\chi^{orb}(j_{!*}P)\geq 0$.
@@
If $g=3$ then $\tau$ is semismall. Therefore $\chi^{orb}({\mathscr{M}}_{3}^{c})\geq 0$
@@
In case (a), $C_{1}$ can be taken to be an elliptic curve with $p$ as its origin. Since $\tau([C])=[\operatorname{Pic}^{0}(C_{1})\times\operatorname{Pic}^{0}(C_{2})]$, the map forgets $p\in C_{2}$. Then we may identify the fibre $\tau^{-1}(\tau([C]))$ with $C_{2}$. In case (b), we take $C_{1}\cap C_{2}$ as the origins of $C_{1}$, $C_{2}$, and $C_{2}\cap C_{3}$ as the origin of $C_{3}$. The point $p\in C_{2}\cap C_{3}$ is distinct from 0 in this case; we treat (c) as the limiting case where $p\to 0$. Then $\tau^{-1}(\tau([C]))$ with $C_{2}$. The claim follows from this, and it implies that $\tau$ is also semismall, when $g=3$. The inequality follows from corollary 4.5. ∎
@@
The preceding proposition now implies our Theorem 1.2 for ${\mathscr{M}}_{3}$ since ${\mathscr{M}}_{3}\hookrightarrow{\mathscr{M}}_{3}^{g}$ is an affine embedding. However, the map ${\mathscr{M}}_{g}^{c}\to{\mathscr{A}}_{g}$ is not semismall when $g>3$, so in general one requires a different strategy. For general ${\mathscr{M}}_{g,n}$ (and their products) we shall use an inductive argument via the partial compactifications ${\mathscr{M}}_{g,n}^{sc}$ (and their products).
As noted in the introduction, we shall prove Theorem 1.6 via induction. The following proposition provides a base case for the induction.
Proposition 7.4.
Let ${\mathscr{M}}={\mathscr{M}}_{0,n_{1}}\times\cdots\times{\mathscr{M}}_{0,n_{r}}\times{}_{N}{\mathscr{A}}_{g_{s},m_{s}}\times\cdots\times{}_{N}{\mathscr{A}}_{g_{1},m_{1}}$, where $g_{i}>0$, $m_{i}\geq 0$, $n_{i}>2$ and $N\geq 3$. If $\mathcal{K}$} is perverse sheaf on $\mathscr{M}$}, then
$\chi({\mathscr{M}},{\mathcal{K}})\geq 0$
Proof.
Write $\mathscr{M}$} as a product $M\times A$ where $M$ is a product of factors of type ${\mathscr{M}}_{0,n}$ and $A$ is a product of factors of type $_{N}{\mathcal{A}}_{g,n}$. Since ${\mathscr{M}}_{0,3}$ is a point, we may assume that $n_{i}>3$. We note that ${\mathscr{M}}_{0,4}\cong{\mathbb{P}}^{1}-\{0,1,\infty\}$. If $E=0+1+\infty$, then $\Omega_{{\mathscr{M}}_{0,4}}^{1}(\log E)$ is nef. Therefore, the proposition holds for any product $({\mathscr{M}}_{0,4})^{l}$ by lemmas 2.2 and 2.3. We have an open immersion $j:{\mathscr{M}}_{0,n}\to({\mathscr{M}}_{0,4})^{n-3}$ sending
$(C,p_{1},\ldots,p_{n})\mapsto((C,p_{1},p_{2},p_{3},p_{i}))_{i=4,\ldots n}$
${\mathscr{M}}_{0,n_{1}}\times\ldots\times{\mathscr{M}}_{0,n_{r}}\xrightarrow{J}({\mathscr{M}}_{0,4})^{\sum n_{i}-3r}=:X$
is also an open affine embedding. By Theorem 4.2, there is a compactification $\overline{A}$ of $A$ such that the log tangent bundle is nef. It follows by Lemma 2.3 that the log cotangent bundle of $X\times\overline{A}$ is nef, and therefore by Lemma 2.2 the conclusion of the proposition holds for $X\times A$. Now $J\times id$ is an open affine embedding of $M\times A$ into $X\times A$. Therefore $(J\times id)_{*}P$ is perverse when $P\in\operatorname{Perv}(M\times A)$. Therefore $\chi^{orb}({\mathscr{M}},P)\geq 0$.
@@
Let ${\mathscr{M}}={\mathscr{M}}_{g_{1},n_{1}}^{sc}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{sc}$, where $2g_{i}-2+n_{i}\geq 3$. If $\mathcal{K}$} is perverse sheaf on $\mathscr{M}$}, then